OpenStudy (anonymous):

For the function y = -2+5sin(pi/12(x-2)), what is the maximum value?

OpenStudy (whpalmer4):

what is the maximum value of 5 sin(x) for any value of x?

OpenStudy (whpalmer4):

add that to -2 and you've got your answer

OpenStudy (anonymous):

x=8+24n ?

OpenStudy (anonymous):

wait what do i do ?

OpenStudy (whpalmer4):

by the way, I'm assuming that is supposed to be interpreted as \[y = -2 + 5\sin(\frac{\pi}{12}(x-2))\]

OpenStudy (anonymous):

yes @whpalmer4

OpenStudy (whpalmer4):

the sin function has a range of -1 to 1, right? so 5*sin has a range of 5*-1 to 5*1 = -5 to 5 So the max value of 5 sin(anything) = 5, and the max value of y will be -2 + 5 =

OpenStudy (anonymous):

3?

OpenStudy (whpalmer4):

is that an answer or a prayer? :-)

OpenStudy (anonymous):

answer :)

OpenStudy (whpalmer4):

okay!

OpenStudy (whpalmer4):

hi @primeralph!

OpenStudy (anonymous):

thank you @primeralph sorry that you felt ignored :(

OpenStudy (anonymous):

I feel bad now :(

OpenStudy (anonymous):

@whpalmer4 so my answer is 3 right? :)

OpenStudy (whpalmer4):

so, because no restrictions were placed on the values of x, we just took the max value produced by sin(x) at any value of x, multiplied by 5, and added in the offset of -2 to get our final answer of 3 being the max. for bonus credit, what is the minimum value of y?

OpenStudy (whpalmer4):

for extra extra credit, what is the first positive value of x that gets us the max value of y :-)

OpenStudy (anonymous):

-2?

OpenStudy (whpalmer4):

nope. the range of sin(x) is -1 to 1, so the range of 5 sin(x) is -5 to 5. take the minimum value (-5) and add to the offset (-2), and you get?

OpenStudy (anonymous):

-7 as the minimum value :)

OpenStudy (whpalmer4):

here's a graph, btw

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

g2g

OpenStudy (whpalmer4):

here's a better one, showing what the sine looks like without the multiplier or offset: